how do I expand my term by using the associative property? (5x+6)/3 = x/5
Thank you,
wildworkz@gmail.com

Answer by rfer(12726)

work this like a proportion
(5x+6)(5)=3x
25x+30=3x
22x=-30
x=-30/22
x=-1 4/11